1. Calculate the pH (give your answers to 2.d.p) of the following strong bases at 298 K.
(a) 0.100 mol dm–3 NaOH
(b) 0.100 mol dm–3 Ca(OH)2
(c) 0.0500 mol dm–3 KOH
1
Expert's answer
2020-11-26T05:08:56-0500

Solution.

i) [OH- ] = 0.100 moldm-3

 [H+] = Kw/[OH- ] = 10-14/0.100 = 10-13  pH = -log[10-13] = 13.00

(ii) [OH- ] = 0.200 moldm-3

 [H+] = Kw/[OH- ] = 10-14/0.100 = 5 x 10-14  pH = -log[5 x 10-14] = 13.30

iii) [OH- ] = 0.050 moldm-3

[H+] = Kw/[OH- ] = 10-14/0.100 = 2 x10-13  pH = -log[2 x 10-13] = 12.69(Ans.)